OpenStudy (anonymous):

Can someone help me with this please? A great runner can run a mile in 4 minutes, which is equivalent to 15 miles per hour. a. If a paramecium can travel 12 lengths of its body in a second, how fast is it going in miles per hour? b. How many times faster than a 4-minute miler can the paramecium move through liquid? c. What factor is ignored in these comparisons? (Hint: it grossly underestimates the paramecium’s speed superiority.)

OpenStudy (tkhunny):

Your basic paramecium comes in \(70\mu m - 140\mu m\) lengths. "a" is just a bad question.

OpenStudy (anonymous):

um how did you come up with that? can you explain please?

OpenStudy (tkhunny):

What's to explain? That's what size they are. Simple as that. Since it varies so much, as much as 100% of the smallest, part "a" needs to my MUCH more specific. Is it 70? Is it 140? What is it?

OpenStudy (anonymous):

oh okay well thank you for your help.. i mean thats the way the problem was presented. I wouldn't know … sorry

OpenStudy (tkhunny):

All i snot lost. After some reality therapy, if we REALLY want to solve the problem, we have to make some assumptions. Why not pick the smallest paramecium. That will be slower than the largest. Okay, then paramecium length is \(70\mu m\). 12 times that is \(840\mu m\). Velocity, then, is \(\dfrac{840\mu m}{1\;sec}\). Convert that to miles / hour.

OpenStudy (anonymous):

I'm sorry I'm lost

OpenStudy (tkhunny):

You can't be lost. We haven't done anything. \(Velocity = \dfrac{Distance}{Time}\). That's all that happened. You may recognize it more as \(Distance = Rate \cdot Time\) -- Same thing.

OpenStudy (anonymous):

OHHHHHH!!! Okay… i got it,… i got!

OpenStudy (anonymous):

thank you

OpenStudy (lena772):

(y)